OpenStudy (anonymous):

The weight of a liquid varies directly as its volume V. If the weight of the liquid in a cubical container 4 cm on a side is 128g, find the weight of the liquid in a cubical container 3cm on a side.

OpenStudy (saifoo.khan):

try it..

OpenStudy (anonymous):

is it 96g

OpenStudy (anonymous):

i dont know if where to start

OpenStudy (saifoo.khan):

How you'd got your answer? :O

OpenStudy (anonymous):

divided 128 by 4 which is 32 then times that by 3

OpenStudy (saifoo.khan):

The answer's correct thought. lol

OpenStudy (anonymous):

wow so i did

OpenStudy (saifoo.khan):

Yes..

OpenStudy (saifoo.khan):

but the steps are, W =kV 128 = k(4) k = 128/4 k = 32 Now, W = 32V W = 32V W = 32(3) W = 96g

OpenStudy (anonymous):

Thanks I am so bad at math

OpenStudy (saifoo.khan):

Dont worry, everyone's the same in the beginning.